OpenStudy (anonymous):

determine whether the expression is a trinomial square w2 - 4w + 4

OpenStudy (dumbcow):

-2*-2 = 4 -2 + -2 = -4 = (w-2)^2

OpenStudy (anonymous):

dumbcow so the ans. would be yes or no

OpenStudy (dumbcow):

well i just factored it into a squared binomial, so its a perfect square

OpenStudy (anonymous):

Ok can yu help with this one factor completely 9r2 + 18rc + 8c2

OpenStudy (dumbcow):

look at factors of 9*8 or 72 12*6 = 72 12 + 6 = 18 (9r +12c)(9r +6c) =(3r+4c)(3r+2c)

OpenStudy (anonymous):

factor completely 2a2 - 28a +98

OpenStudy (dumbcow):

do the same steps, look at factors of 2*98 or 196

OpenStudy (dumbcow):

or actually divide everything by 2 first 2(a^2 -14a +49)

OpenStudy (dumbcow):

look at factors of 49, what adds up to -14

OpenStudy (anonymous):

I came up with this answer is it correct 2(a-7)2

OpenStudy (dumbcow):

yes, good job

OpenStudy (anonymous):

thnks for the clarification

OpenStudy (anonymous):

factor the trinomial 20y2 + 24y - 32 Can u check my answer pls: 4(5y-4)(y+2)

OpenStudy (dumbcow):

correct :)

OpenStudy (anonymous):

thnks

OpenStudy (anonymous):

can you check my answer plse: factor 42+45y-27v2 my ans is: -3(3v+2)(3v-7)

OpenStudy (dumbcow):

correct